# Single-photon scattering in giant-atom topological-waveguide-QED systems

### v1

@@ Abstract
The giant-atom topological-waveguide-QED systems have recently emerged as a promising platform for manipulating light-matter interactions. The combination of the multiple-point couplings and topological phase effect could lead to rich physical phenomena and effects. Here, we study single-photon scattering in a Su-Schrieffer-Heeger (SSH) waveguide coupled to either one or two two-level giant atoms. We assume that each giant atom is coupled to the waveguide via two coupling points, and hence there exist four and sixteen coupling configurations for the single-giant-atom case and two-giant-atom separate coupling case, respectively. By solving the single-photon scattering problem in the real space, we obtain the exact expressions of the single-photon scattering amplitudes. It is found that a single photon in the SSH waveguide can be completely reflected or transmitted by choosing proper coupling configurations, coupling-point distances, atomic resonance frequency, and dimerization parameter. In addition, under proper parameter conditions, the scattering spectra are periodically modulated by the coupling-point distances. We also find that the collective behavior of the two giant atoms can be adjusted by quantum interference effect and topological effect, and that the single-photon scattering spectra can exhibit the Lorentzian, super-Gaussian, electromagnetically induced transparency-like, and asymmetric Fano line shapes for some coupling configurations. This work will inspire the development of controllable single-photon devices based on the giant-atom topological-waveguide-QED systems.
@@ I Introduction
The waveguide quantum electrodynamics (QED) systems Sheremet2023 ; Roy2017 ; Gu2017 , as an excellent platform for exploring quantum light-matter interactions, have attracted increasing interest due to their potential applications in quantum optics Blais1997 ; Tudela2024 and quantum information science Nielsen2000 ; Zheng2013 . Many interesting physical effects, such as bound state John1990 ; Shen2007 ; Liao2010a ; Liao2010b ; Hood2016 ; Shi2016 ; Burillo2017 ; Fong2017 ; Calajo2019 ; Mahmoodian2020 , super/subradiance vanLoo2013 ; Zanner2022 , quantum routers Hoi2011 ; Zhou2013 ; Shomroni2014 , and few-photon scattering Shen2005a ; Shen2005b ; Chang2007 ; Zhou2008 ; Tsoi2008 ; Liao2009 ; Shi2009 ; Tsoi2009 ; Liao2013 ; Neumeier2013 ; Liao2016 ; Ke2019 have been widely investigated in waveguide-QED systems. In particular, the waveguide-QED systems can serve good platforms to transfer quantum information and have potential applications in realizing large-scale quantum networks.
@@
In recent years, the giant-atom waveguide-QED systems Kockum2014 ; Kockum125 ; Kannan2020 have attracted extensive attention from the peers in the fields of quantum optics and quantum information. In most quantum optics systems, atoms are typically treated as point-like objects Walls2008 ; Scully1997 . This is because the size of natural atoms is much smaller than the wavelengths of the coupled electromagnetic fields. Differently, the size of the giant atoms is comparable to or even larger than the wavelength of the coupled fields. Therefore, the giant atoms can no longer be considered as point-like objects, and then the dipole approximation becomes invalid. Instead, the giant atoms interact with the electromagnetic fields in the waveguide at multiple coupling points, and the path quantum interferences become more abundant. So far, many interesting physical phenomena have been found in giant-atom waveguide-QED systems, including frequency-dependent Lamb shift and dissipation rate Kockum2014 ; Vadiraj2021 ; Yu2021 ; ZQWang2022 ; Joshi2023 , decoherence-free interaction between giant atoms Kannan2020 ; Kockum2018 ; Carollo2020 ; Soro2023 , non-Markovian dynamics Guo2017 ; Andersson2019 ; Du2021 ; Yin2022a ; Qiu2023 , single-photon scattering Zhao2020 ; Cai2021 ; Feng2021 ; Yin2022b ; Zhu2022 ; Xiang2022 , and chiral light-matter interactions XWang2021 ; Du2022 ; Soro2022 ; XWang2022 ; Zhou2023 .
@@
In parallel, much recent attention in waveguide-QED has been paid to the interaction of atoms with topological waveguides Hasan2010 ; Wra2010 ; Barik2018 ; Ozawa2019 ; Shalaev2019 . Some topics in the small-atom topological-waveguide-QED systems have been studied Dusuel2011 ; Tudela2015 ; YWang2020 ; LJWang2020 ; Kim2021 ; Dong2021 ; Leonforte2021 ; Dong2022 ; Hauff2022 ; Vega2023 , such as single-photon scattering Barik2018 ; LJWang2020 ; Kim2021 ; Bello2019 and topologically protected quantum entanglement Rechtsman2016 ; Mittal2019 ; Wang2019 . Meanwhile, people studied the physical phenomena related to the topological effect of waveguide and the interference effect of giant atoms in giant-atom topological-waveguide-QED systems Vega2021 ; Cheng2022 ; Bag2023 ; Luo2024 ; Wang2024 . Nevertheless, how does the topological effect affect the interference effect remain unclear. In addition, when considering the giant atoms coupled to a topological waveguide, the combined effects of both the multiple path quantum interference for giant atoms and the topological phase of the waveguide on the single-photon scattering behavior is an interesting topic to be investigated.
@@
In this paper, we study the single-photon scattering in a photonic Su-Schrieffer-Heeger (SSH) waveguide Su1979 coupled to either one or two two-level giant atoms. By solving the probability amplitude equations in real space, we obtain the exact expressions of the single-photon scattering amplitudes for the two cases. It is shown that the single-photon scattering behavior depends on the coupling configurations, coupling-point distances, atomic resonance frequency, and dimerization parameter. In the weak-coupling regime, the scattering spectra can exhibit the Lorentzian, super-Gaussian, electromagnetically induced transparency (EIT)-like Abi-Salloum2010 ; Anisimov2011 , and asymmetric Fano line shapes Fano1961 ; Miroshnichenko2010 . In particular, by adjusting the quantum interference and topological effects, the incident photon can be transmitted completely. In the single-giant-atom case, we analyze in detail the influence of the system parameters on the scattering spectra. We find that the scattering spectra are periodically modulated by the coupling-point distance. In the two-giant-atom case, we discuss the conditions for the appearance of different line shapes. In particular, we show that when any two or more legs of the giant atoms are coupled to the sublattices of different types, the corresponding scattering spectra depend on the topological characteristics of the system.
@@ II.1 Physical model and Hamiltonian
We consider a single two-level giant-atom topological-waveguide-QED system, as shown in Fig. 1(a). Here, the giant atom is coupled to a one-dimensional discrete topological waveguide through two separate coupling points, and the waveguide is formed by a photonic SSH-type coupled cavity array. The total Hamiltonian of this system reads
@@
where $\omega_{e}$ is the energy separation between the excited state $\left|e\right\rangle$ and the ground state $\left|g\right\rangle$ of the giant atom. In this work, we set the energy of the ground state $\left|g\right\rangle$ as zero. The $\hat{\sigma}^{+}=\left|e\right\rangle\left\langle g\right|$ and $\hat{\sigma}^{-}=\left|g\right\rangle\left\langle e\right|$ are, respectively, the rising and lowing operators of the giant atom.
@@
where $\hat{a}_{j}^{\dagger}$ $(\hat{b}_{j}^{\dagger})$ and $\hat{a}_{j}$ $(\hat{b}_{j})$ are, respectively, the creation and annihilation operators of photons at sublattice A (B) in the $j$th unit cell. The intracell and intercell hopping rates are $\xi_{1}=J\left(1+\delta\right)$ and $\xi_{2}=J\left(1-\delta\right)$, respectively, with $\delta$ being the dimerization parameter. For $\delta>0$ ($\delta<0$), the SSH waveguide is in its trivial (topological) phase Asboth2016 . To calculate the dispersion relation of the SSH waveguide, we consider the periodic boundary conditions and perform the discrete Fourier transforms
@@
For simplicity, we take the photonic lattice constant $d_{0}=1$. Here, the range of the wave vector $k$ in the first Brillouin zone is $k\in[-\pi,\pi)$.
The Hamiltonian of the SSH waveguide can be transformed into $\hat{H}_{wg}=\sum_{k}\hat{Q}_{wg}^{\dagger}h_{wg}\hat{Q}_{wg}$, where we introduce $\hat{Q}_{wg}=(\hat{a}_{k},\hat{b}_{k})^{T}$ and the kernel
$h_{wg}=\left(\begin{array}[]{cc}\omega_{a}&y\left(k\right)\\ y^{\ast}\left(k\right)&\omega_{a}\end{array}\right),$ (7)
@@
The single giant atom is coupled to the topological waveguide via multiple coupling points. Here, each coupling point is still described by the atom-field interaction under the dipole approximation. We consider the weak-coupling case, then the interaction Hamiltonian can be written under the rotating-wave approximation (RWA) as
@@ II.2 Scattering probability amplitudes
where $\Delta=\omega_{e}-\omega_{a}$ is the frequency detuning between the giant atom and the cavity field. The total excitation number operator $\hat{N}=\hat{\sigma}^{+}\hat{\sigma}^{-}+\sum_{j=1}^{N}(\hat{a}_{j}^{\dagger}\hat{a}_{j}+\hat{b}_{j}^{\dagger}\hat{b}_{j})$ of the system is a conserved quantity, and thus in the single-excitation subspace, a general state of the system can be written as
@@
The probability amplitudes can be obtained by solving the stationary Schrödinger equation $\hat{H}_{\text{rot},1}\left|\Psi_{k}\right\rangle_{1}=\omega^{(\pm)}\left|\Psi_{k}\right\rangle_{1}$. To this end, we assume that a single photon with energy $\omega^{(\pm)}=\pm\omega_{k}$ is initially injected from the left-hand side of the waveguide. For the cases of $j<n$, $n<j<m$, and $j>m$, the probability amplitudes $u_{k}(j)$ and $w_{k}(j)$ corresponding to the upper and lower bands take the form conditions
@@
In Eqs. (16d) and (16h), $r_{1}$ and $t_{1}$ are single-photon reflection and transmission amplitudes, respectively. The variables $A$ and $B$ are the probability amplitudes related to the right- and left-propagating fields between the two coupling points of giant atom. By using the continuity conditions at the $n$th and $m$th coupling points, we have the relations
@@ II.3 Scattering spectra
For convenience, we introduce the accumulated phases $\varphi=\varphi^{\text{AA}}=\varphi^{\text{BB}}=kd$, $\varphi^{\text{AB}}=kd-\phi_{k}$, and $\varphi^{\text{BA}}=kd+\phi_{k}$ of photons propagating between the two coupling points. Here, the superscript represents the different coupling configurations. For the AB- and BA-coupling cases, the phases $\varphi^{\text{AB}}$ and $\varphi^{\text{BA}}$ satisfy the relation
@@
Figure 2: The wave vector $k_{\Delta}$ in Eq. (21) as a function of the frequency detuning $\Delta$ and the dimerization parameter $\delta$. The black dashed curve is used to label $k_{\Delta}=-\pi/2$. The area enclosed by the white dashed line determines the range of $\Delta$ and $\delta$.
Base on the above analyses, below we focus on the single-photon scattering in the AA- and AB-coupling cases. Concretely, we analyze the influence of the coupling configurations, coupling-point distances, frequency detuning, and dimerization parameter on the single-photon scattering. We assume that the incident photon resonates with the giant atom, i.e., $\omega_{k}=\Delta$, then the wave vector $k$ can be expressed as
@@
AA $\Delta_{L,k_{\Delta}}^{\text{AB}}$ 0.5 $-\Gamma_{e}\sin\phi_{k_{\Delta},+}$ $-0.5$ $2\Gamma_{e}(1+\sin\phi_{k_{\Delta},+})$
$2\Gamma_{e}(1-\cos\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1-\sin\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1+\cos\phi_{k_{\Delta},+})$
Figure 3: Reflection coefficient $R_{1}$ as a function of the detuning $\Delta_{k}$ at the trivial phase ($\delta>0$) and the topological phase ($\delta<0$) for (a) and (b) the AA-coupling configuration, (c) and (d) the AB-coupling configuration. The left and right columns correspond to the odd and even coupling-point distances, respectively. Other parameters used are $g=0.01J$ and $\Delta\approx 1.58J$ ($k_{\Delta}=-\pi/2$).
We firstly study the influence of the sign of the dimerization parameter $\delta$ and the coupling-point distance $d$ on the single-photon scattering. Specifically, we take $\delta=\pm 0.5$ and $k_{\Delta}=-\pi/2$. According to Eq. (21), the corresponding frequency detuning is $\Delta\approx 1.58J$. In Fig. 3, we plot the reflection coefficient $R_{1}$ for the AA- and AB-coupling configurations as a function of the detuning $\Delta_{k}$ when $\delta$ and $d$ take different values. For the AA-coupling configuration, the sign of $\delta$ has no influence on the reflection spectra, which is consistent with the previous analyses. However, the locations of the reflection peaks can be adjusted by changing the values of $d$. In particular, when $d=1$ and $3$, we find that the reflection spectra are the Lorentzian line shapes, as shown in Fig. 3(a). We also find that the two Lorentzian line shapes are symmetric with respect to $\Delta_{k}=0$ and have equal linewidths. When $d=2$ and $4$, the reflection spectra in Fig. 3(b) exhibit remarkably different features. For $d=4$, the reflection spectrum exhibits the Lorentzian line shape with reflection peak located at $\Delta_{k}=0$. Differently, for $d=2$, the incident photon is completely transmitted ($R_{1}=0$). This phenomenon is caused by the constructive interference between the two coupling channels of the giant atom Kockum125 .
For the AB-coupling configuration, we find that the reflection spectra are characterized by the Lorentzian line shapes, as shown in Figs. 3(c) and 3(d). Different from the AA-coupling configuration, both the sign of $\delta$ and the value of $d$ can affect the reflection spectra of the AB-coupling configuration. Though the reflection spectra are the Lorentzian line shapes for both $\delta=0.5$ and $-0.5$, the sign of $\delta$ can change the reflection peak and linewidth. This means that the single-photon scattering for the AB-coupling configuration is topologically dependent. Moreover, we find that the red solid curve in Fig. 3(c) (with $\delta=0.5$ and $d=1$) and the red dashed curve with dots in Fig. 3(d) (with $\delta=-0.5$ and $d=2$) are symmetric with respect to $\Delta_{k}=0$, and vice versa. These two reflection spectra have equal linewidths. The same relationship is observed for the reflection spectra for $d=3$ and $d=4$ [see the blue solid curve and dashed curve with dots in Figs. 3(c) and 3(d)].
To better explain these phenomena, we only consider the case where the incident photons in the vicinity of atomic transition frequency can effectively interact with the atom. Consequently, we can replace the wave vector $k$ in Eqs. (19a)-(LABEL:phase) with $k_{\Delta}$. In this case, the characteristic quantities in Eqs. (19a)-(LABEL:phase) for the upper-band case can be approximated as
@@
where the corresponding propagating phases are $\varphi_{k_{\Delta}}=k_{\Delta}d$, $\varphi_{k_{\Delta}}^{\text{AB}}=k_{\Delta}d-\phi_{k_{\Delta}}$, and $\varphi_{k_{\Delta}}^{\text{BA}}=k_{\Delta}d+\phi_{k_{\Delta}}$. According to Eqs. (22a)-(LABEL:approximated:phase), Eq. (18a) can be approximated as

### v2

@@ Abstract
The giant-atom topological-waveguide-QED systems have recently emerged as a promising platform for manipulating light-matter interactions. The combination of the multiple-point couplings and topological phase effect could lead to rich physical phenomena and effects. Here, we study single-photon scattering in a Su-Schrieffer-Heeger (SSH) waveguide coupled to either one or two two-level giant atoms. We assume that each giant atom is coupled to the waveguide via two coupling points and hence there exist four and sixteen coupling configurations for the single-giant-atom case and two-giant-atom separate coupling case, respectively. By solving the single-photon scattering problem in the real space, we obtain the exact expressions of the single-photon scattering amplitudes. It is found that a single photon in the SSH waveguide can be completely reflected or transmitted by choosing proper coupling configurations, coupling-point distances, atomic resonance frequency, and dimerization parameter. In addition, under proper parameter conditions, the scattering spectra are periodically modulated by the coupling-point distances. We also find that the collective behavior of the two giant atoms can be adjusted by quantum interference effect and topological effect and that the single-photon scattering spectra can exhibit the Lorentzian, super-Gaussian, electromagnetically induced transparencylike, and asymmetric Fano line shapes for some coupling configurations. This work will inspire the development of controllable single-photon devices based on the giant-atom topological-waveguide-QED systems.
@@ I Introduction
The waveguide quantum electrodynamics (QED) systems Sheremet2023 ; Roy2017 ; Gu2017 , as an excellent platform for exploring quantum light-matter interactions, have attracted increasing interest due to their potential applications in quantum optics Blais1997 ; Tudela2024 and quantum information science Nielsen2000 ; Zheng2013 . Many interesting physical effects, such as bound state John1990 ; Shen2007 ; Liao2010a ; Liao2010b ; Hood2016 ; Shi2016 ; Burillo2017 ; Fong2017 ; Calajo2019 ; Mahmoodian2020 , super- and subradiance vanLoo2013 ; Zanner2022 , quantum routers Hoi2011 ; Zhou2013 ; Shomroni2014 , and few-photon scattering Shen2005a ; Shen2005b ; Chang2007 ; Zhou2008 ; Tsoi2008 ; Liao2009 ; Shi2009 ; Tsoi2009 ; Liao2013 ; Neumeier2013 ; Liao2016 ; Ke2019 have been widely investigated in waveguide-QED systems. In particular, the waveguide-QED systems can serve as good platforms to transfer quantum information and have potential applications in realizing large-scale quantum networks.
@@
In recent years, the giant-atom waveguide-QED systems Kockum2014 ; Kockum125 ; Kannan2020 have attracted extensive attention from the peers in the fields of quantum optics and quantum information. In most quantum optics systems, atoms are typically treated as pointlike objects Walls2008 ; Scully1997 . This is because the size of natural atoms is much smaller than the wavelengths of the coupled electromagnetic fields. Differently, the size of the giant atoms is comparable to or even larger than the wavelength of the coupled fields. Therefore, the giant atoms can no longer be considered as pointlike objects and then the dipole approximation becomes invalid. Instead, the giant atoms interact with the electromagnetic fields in the waveguide at multiple coupling points and the path quantum interferences become more abundant. So far, many interesting physical phenomena have been found in giant-atom waveguide-QED systems, including frequency-dependent Lamb shift and dissipation rate Kockum2014 ; Vadiraj2021 ; Yu2021 ; ZQWang2022 ; Joshi2023 , decoherence-free interaction between giant atoms Kannan2020 ; Kockum2018 ; Carollo2020 ; Soro2023 , non-Markovian dynamics Guo2017 ; Andersson2019 ; Du2021 ; Yin2022a ; Qiu2023 , single-photon scattering Zhao2020 ; Cai2021 ; Feng2021 ; Yin2022b ; Zhu2022 ; Xiang2022 ; Zhao2024 , and chiral light-matter interactions XWang2021 ; Du2022 ; Soro2022 ; XWang2022 ; Zhou2023 ; Roccati2024L . These studies consider the coupling between the giant atoms and either linear or tight-binding waveguides, focusing on the multipoint interference effects. Compared to the linear waveguide, the tight-binding waveguide provides a highly structured environment for the giant atoms and allows them to couple to multiple resonators Zhao2020 . In addition, photons propagating in the tight-binding waveguide exhibit a nonlinear dispersion relation and the photonic group velocity can be tuned Zhou2008 ; Zhao2020 .
@@
In parallel, much recent attention in waveguide-QED has been paid to the interaction of atoms with topological waveguides Hasan2010 ; Wra2010 ; Barik2018 ; Ozawa2019 ; Shalaev2019 . Different from the linear and tight-binding waveguides, the SSH waveguide not only provides a highly structured coupling environment for the giant atoms, but also exhibits topological characteristics. Some topics in the small-atom topological-waveguide-QED systems have been studied Dusuel2011 ; Tudela2015 ; YWang2020 ; LJWang2020 ; Kim2021 ; Dong2021 ; Leonforte2021 ; Dong2022 ; Hauff2022 ; Vega2023 ; Roccati2024NC ; Lu2024 , such as single-photon scattering Barik2018 ; LJWang2020 ; Kim2021 ; Bello2019 and topologically protected quantum entanglement Rechtsman2016 ; Mittal2019 ; Wang2019 ; Dai2022 . Meanwhile, people studied the physical phenomena related to the topological effect of waveguide and the interference effect of giant atoms in giant-atom topological-waveguide-QED systems Vega2021 ; Cheng2022 ; Bag2023 ; JJWang2023 ; Luo2024 ; Wang2024 . Nevertheless, how the topological effect influences the interference effect remains unclear and the various coupling configurations in giant-atom topological waveguide-QED systems are also intriguing. Therefore, when considering the giant atoms coupled to a topological waveguide, the combined effects of both the multipath quantum interference for giant atoms and the topological phase of the waveguide on the single-photon scattering behavior is an interesting topic to be investigated.
@@
In this paper, we study the single-photon scattering in a photonic Su-Schrieffer-Heeger (SSH) waveguide Su1979 coupled to either one or two two-level giant atoms. In this case, the incident photon in the SSH waveguide not only has an adjustable group velocity, but also will be influenced by the topological characteristics of the waveguide. By solving the probability amplitude equations in real space, we obtain the exact expressions of the single-photon scattering amplitudes for the two cases. We also find that the scattering amplitudes are determined by the Lamb shift, exchanging interaction, individual decay, and collective decay. The phases of these quantities depend on either the coupling-point distances or the dimerization parameter. It is shown that the single-photon scattering behavior depends on the coupling configurations, coupling-point distances, atomic resonance frequency, and dimerization parameter. In the weak-coupling regime, the scattering spectra can exhibit the Lorentzian, super-Gaussian, electromagnetically induced transparency (EIT)-like Abi-Salloum2010 ; Anisimov2011 , and asymmetric Fano line shapes Fano1961 ; Miroshnichenko2010 . In particular, by adjusting the quantum interference and topological effects, the incident photon can be transmitted completely. In the single-giant-atom case, we analyze in detail the influence of the system parameters on the scattering spectra. We find that the scattering spectra are periodically modulated by the coupling-point distance. In the two-giant-atom case, we discuss the conditions for the appearance of different line shapes and the relationships between the reflection coefficients for different coupling configurations. In particular, we show that when any two or more legs of the giant atoms are coupled to the sublattices of different types, the phases in the scattering amplitudes include the topology-dependent phase. Consequently, the corresponding scattering spectra depend on the topological characteristics of the system.
@@ II.1 Physical model and Hamiltonian
We consider a single two-level giant-atom topological-waveguide-QED system, as shown in Fig. 1(a). Here, the giant atom is coupled to a one-dimensional discrete topological waveguide through two separate coupling points and the waveguide is formed by a photonic SSH-type coupled cavity array. The total Hamiltonian of this system reads
@@
where $\omega_{e}$ is the energy separation between the excited state $\left|e\right\rangle$ and the ground state $\left|g\right\rangle$ of the giant atom. In this work, we set the energy of the ground state $\left|g\right\rangle$ as zero. The $\hat{\sigma}^{+}=\left|e\right\rangle\left\langle g\right|$ and $\hat{\sigma}^{-}=\left|g\right\rangle\left\langle e\right|$ are, respectively, the raising and lowering operators of the giant atom.
@@
where $\hat{a}_{j}^{\dagger}$ $(\hat{b}_{j}^{\dagger})$ and $\hat{a}_{j}$ $(\hat{b}_{j})$ are, respectively, the creation and annihilation operators of photons at sublattice A (B) in the $j$th ($j=1,2,3,...,N$) unit cell. The intracell and intercell hopping rates are $\xi_{1}=J\left(1+\delta\right)$ and $\xi_{2}=J\left(1-\delta\right)$, respectively, with $\delta$ being the dimerization parameter. For $\delta>0$ ($\delta<0$), the SSH waveguide is in its trivial (topological) phase Asboth2016 . To calculate the dispersion relation of the SSH waveguide, we consider the periodic boundary conditions and perform the discrete Fourier transforms
@@
For simplicity, we take the photonic lattice constant $d_{0}=1$. Here, the wave vector in the first Brillouin zone is $k\in[-\pi,...,\pi-2\pi/N]$.
The Hamiltonian of the SSH waveguide can be transformed into $\hat{H}_{wg}=\sum_{k}\hat{Q}_{k}^{\dagger}h_{k}\hat{Q}_{k}$, where we introduce $\hat{Q}_{k}=(\hat{a}_{k},\hat{b}_{k})^{T}$ and the kernel
$h_{k}=\left(\begin{array}[]{cc}\omega_{a}&y\left(k\right)\\ y^{\ast}\left(k\right)&\omega_{a}\end{array}\right),$ (7)
@@
The single giant atom is coupled to the topological waveguide via multiple coupling points. Here, each coupling point is still described by the atom-field interaction under the dipole approximation. We consider the weak-coupling case; then the interaction Hamiltonian can be written under the rotating-wave approximation (RWA) as
@@ II.2 Scattering probability amplitudes
where $\Delta=\omega_{e}-\omega_{a}$ is the frequency detuning between the giant atom and the cavity field. The total excitation number operator $\hat{N}=\hat{\sigma}^{+}\hat{\sigma}^{-}+\sum_{j=1}^{N}(\hat{a}_{j}^{\dagger}\hat{a}_{j}+\hat{b}_{j}^{\dagger}\hat{b}_{j})$ of the system is a conserved quantity and thus, in the single-excitation subspace, a general state of the system can be written as
@@
The probability amplitudes can be obtained by solving the stationary Schrödinger equation $\hat{H}_{\text{rot},1}\left|\Psi_{k}\right\rangle_{1}=\omega^{(\pm)}\left|\Psi_{k}\right\rangle_{1}$. To this end, we consider that a single photon with energy $\omega^{(\pm)}=\pm\omega_{k}$ is initially injected from the left-hand side of the waveguide. For the cases of $j<n$, $n<j<m$, and $j>m$, the probability amplitudes $u_{k}(j)$ and $w_{k}(j)$ corresponding to the upper and lower bands take the form conditions
@@
In Eqs. (16d) and (16h), $r_{1}$ and $t_{1}$ are single-photon reflection and transmission amplitudes, respectively. The variables $A$ and $B$ are the probability amplitudes related to the right- and left-propagating fields between the two coupling points of the giant atom. By using the continuity conditions at the $n$th and $m$th coupling points, we have the relations
@@ II.3 Scattering spectra
For convenience, we introduce the accumulated phases $\varphi=\varphi^{\text{AA}}=\varphi^{\text{BB}}=kd$, $\varphi^{\text{AB}}=kd-\phi_{k}$, and $\varphi^{\text{BA}}=kd+\phi_{k}$ of photons propagating between the two coupling points. Here, the superscript represents the different coupling configurations. We find that, for the AA- and BB-coupling cases, the phase is only related to the coupling-point distance $d$, while for the AB- and BA coupling-cases, the phases $\varphi^{\text{AB}}$ and $\varphi^{\text{BA}}$ also include the topology-dependent phase $\phi_{k}$. Therefore, the single-photon scattering behavior for these two coupling configurations is highly sensitive to the topological characteristics of the system. Moreover, we find that the phases $\varphi^{\text{AB}}$ and $\varphi^{\text{BA}}$ satisfy the relation
@@
Based on the above analyses, below we focus on the single-photon scattering in the AA- and AB-coupling cases. Concretely, we analyze the influence of the coupling configurations, coupling-point distances, frequency detuning, and dimerization parameter on the single-photon scattering. Note that the atomic resonance frequency $\Delta$ can lie within the band gap, inside the energy bands, or outside the energy bands. Some previous studies of the giant atom-topological waveguide QED systems have shown that atom-photon bound states can appear in the system when the atomic resonance frequency lies within the band gap or outside the bands Bello2019 ; Vega2021 ; Cheng2022 ; Wang2024 . In particular, when the atomic resonance frequency lies within the band gap (for example $\Delta=0$), the giant atom can act as an effective boundary and induce the chiral zero-energy modes (i.e., the chiral bound states) in the waveguide under the periodic boundary condition Cheng2022 . The properties of these chiral bound states are similar to the edge states under the open-boundary SSH waveguide, but their formation conditions and distributions are different Bello2019 ; Vega2021 ; Cheng2022 ; Wang2024 . First, the chiral bound states and the edge states correspond to periodic and open boundary conditions of the waveguide, respectively. Second, the photonic probability distributions of chiral bound states are located at the left side of the giant atom or between the two coupling points for A-A coupling Cheng2022 . The edge states are mostly located at the two ends of the SSH waveguide Cheng2022 ; Lang2012 . Here, we focus on the single-photon scattering with the incident-photon frequency within the energy band. It is found that when the atomic resonance frequency lies within the band gap or outside the energy band, the incident photon is completely transmitted. In contrast, when the atomic resonance frequency is within the energy band and resonates with the incident photon ($\omega_{k}=\Delta$), the photon can be scattered. In this case, the wave vector $k$ can be expressed as
@@
Figure 2: Wave vector $k_{\Delta}$ in Eq. (21) as a function of the frequency detuning $\Delta$ and the dimerization parameter $\delta$. The black dashed curve is used to label $k_{\Delta}=-\pi/2$. The area enclosed by the white dashed line determines the range of $\Delta$ and $\delta$.
We first study the influence of the sign of the dimerization parameter $\delta$ and the coupling-point distance $d$ on the single-photon scattering. Specifically, we take $\delta=\pm 0.5$ and $k_{\Delta}=-\pi/2$. According to Eq. (21), the corresponding frequency detuning is $\Delta\approx 1.58J$. In Fig. 3, we plot the reflection coefficient $R_{1}$ for the AA- and AB-coupling configurations as a function of the detuning $\Delta_{k}$ when $\delta$ and $d$ take different values. For the AA-coupling configuration, the sign of $\delta$ has no influence on the reflection spectra, which is consistent with the previous analyses. However, the locations of the reflection peaks can be adjusted by changing the values of $d$. In particular, when $d=1$ and $3$, we find that the reflection spectra are the Lorentzian line shapes, as shown in Fig. 3(a). We also find that the two Lorentzian line shapes are symmetric with respect to $\Delta_{k}=0$ and have equal linewidths. When $d=2$ and $4$, the reflection spectra in Fig. 3(b) exhibit remarkably different features. For $d=4$, the reflection spectrum exhibits the Lorentzian line shape with reflection peak located at $\Delta_{k}=0$. Differently, for $d=2$, the incident photon is completely transmitted ($R_{1}=0$). This phenomenon is caused by the constructive interference between the two coupling channels of the giant atom Kockum125 .
Figure 3: Reflection coefficient $R_{1}$ as a function of the detuning $\Delta_{k}$ at the trivial phase ($\delta>0$) and the topological phase ($\delta<0$) for (a) and (b) the AA-coupling configuration and (c) and (d) the AB-coupling configuration. The left and right columns correspond to the odd and even coupling-point distances, respectively. Other parameters used are $g=0.01J$ and $\Delta\approx 1.58J$ ($k_{\Delta}=-\pi/2$).
For the AB-coupling configuration, we find that the reflection spectra are characterized by the Lorentzian line shapes, as shown in Figs. 3(c) and 3(d). Different from the AA-coupling configuration, both the sign of $\delta$ and the value of $d$ can affect the reflection spectra of the AB-coupling configuration. Though the reflection spectra are the Lorentzian line shapes for both $\delta=0.5$ and $-0.5$, the sign of $\delta$ can change the reflection peak and linewidth. This means that the single-photon scattering for the AB-coupling configuration is topologically dependent. Moreover, we find that the red solid curve in Fig. 3(c) (with $\delta=0.5$ and $d=1$) and the red dashed curve with dots in Fig. 3(d) (with $\delta=-0.5$ and $d=2$) are symmetric with respect to $\Delta_{k}=0$ and vice versa. These two reflection spectra have equal linewidths. The same relationship is observed for the reflection spectra for $d=3$ and $d=4$ [see the blue solid curve and dashed curve with dots in Figs. 3(c) and 3(d)]. We should mention that the dependence of the single-photon scattering on the topological feature of the waveguide can also be studied by examining the dependence of the scattering behavior on the Zak phase. A relating experimental study has been realized in periodic acoustic systems Xiao2015 . The relation between the scattering feature and the Zak phase can be directly obtained based on the correspondence between $\delta>0$ ($\delta<0$) and the 00 ($\pi$) Zak phase.
@@
AA $\Delta_{L,k_{\Delta}}^{\text{AA}}$$\pm 0.5$ $-\Gamma_{e}$ 00 $\Gamma_{e}$ 00
$\Gamma_{k_{\Delta}}^{\text{AA}}$$\pm 0.5$ $2\Gamma_{e}$ 00 $2\Gamma_{e}$ $4\Gamma_{e}$
AB $\Delta_{L,k_{\Delta}}^{\text{AB}}$ 0.5 $-\Gamma_{e}\sin\phi_{k_{\Delta},+}$ $-\Gamma_{e}\cos\phi_{k_{\Delta},+}$ $\Gamma_{e}\sin\phi_{k_{\Delta},+}$ $\Gamma_{e}\cos\phi_{k_{\Delta},+}$ $-\Gamma_{e}\sin\phi_{k_{\Delta},+}$
$-0.5$ $-\Gamma_{e}\cos\phi_{k_{\Delta},+}$ $-\Gamma_{e}\sin\phi_{k_{\Delta},+}$ $\Gamma_{e}\cos\phi_{k_{\Delta},+}$ $\Gamma_{e}\sin\phi_{k_{\Delta},+}$ $-\Gamma_{e}\cos\phi_{k_{\Delta},+}$
$\Gamma_{k_{\Delta}}^{\text{AB}}$ 0.5 $2\Gamma_{e}(1+\cos\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1-\sin\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1-\cos\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1+\sin\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1+\cos\phi_{k_{\Delta},+})$
$-0.5$ $2\Gamma_{e}(1+\sin\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1-\cos\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1-\sin\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1+\cos\phi_{k_{\Delta},+})$ $2\Gamma_{e}(1+\sin\phi_{k_{\Delta},+})$
To better explain these phenomena, we only consider the case where the incident photons in the vicinity of atomic transition frequency can effectively interact with the atom. Consequently, we can replace the wave vector $k$ in Eqs. (19a)–(LABEL:phase) with $k_{\Delta}$. In this case, the characteristic quantities in Eqs. (19a)–(LABEL:phase) for the upper-band case can be approximated as
@@
where the corresponding propagating phases are $\varphi_{k_{\Delta}}=k_{\Delta}d$, $\varphi_{k_{\Delta}}^{\text{AB}}=k_{\Delta}d-\phi_{k_{\Delta}}$, and $\varphi_{k_{\Delta}}^{\text{BA}}=k_{\Delta}d+\phi_{k_{\Delta}}$. According to Eqs. (22a)–(LABEL:approximated:phase), Eq. (18a) can be approximated as